# Foundation Models on a Budget: Approximating Blocks in Large Vision Models

### v5

@@ 6 Conclusion
In this work, we first analyze the emergence of consistent block-wise representation similarities within pretrained foundation models and then propose a method to leverage these similarities to obtain smaller yet performant models. To this end, we propose Transformer Blocks Approximation (TBA), a novel method for identifying and efficiently approximating similar transformer blocks using a simple linear transformation, without requiring additional training or fine-tuning. Our extensive empirical evaluations across multiple pretrained vision models and datasets validate that TBA significantly reduces model parameters while maintaining, and sometimes even improving, downstream task performance. Notably, the learned linear transformations exhibit strong generalization capabilities across datasets, enabling lightweight model adaptation. Furthermore, TBA’s straightforward linear approach often achieves better results than existing strategies like block skipping, and can be more effective than complex, trained approximations. TBA thus offers a practical and efficient method for streamlining foundation models, making them more computationally accessible.
@@ A.1 Implementation details
This section details the experiments conducted in Section 4, providing information to reproduce them. Additionally, we provide the code as a zip file in the supplementary material.
@@ A.1.1 Models and Datasets
Table 6 contains the full list of the pretrained models, while Table 7 contains dataset information.
@@ A.2 Additional Experiments
This section presents supplementary experiments to extend those detailed in Section 4.
A.2.1 Block Similarities
The analysis in Section 4.1 reveals distinct block-wise similarity patterns within pretrained foundation models. The analysis utilizes the mean of token representations within each block to compute similarities. However, as detailed in Figure 5, these characteristic similarity patterns persist and are qualitatively very similar when using only the CLS token’s representation from each block. This suggests a degree of interchangeability for this type of analysis.
Finally, to assess the impact of the similarity measure itself, we conducted an ablation study correlating various metrics with downstream task performance on ViT-S and ImageNet1k (Figure 6). Our findings reveal that even simple metrics, such as MSE or cosine similarity, when used to identify block-level similarities, exhibit a strong correlation with final accuracy. These results underscore the robustness and flexibility of leveraging such block similarities, indicating compatibility with a range of similarity or distance functions capable of capturing these inter-block relationships.
| Approx. | Params. | ImageNet1k |
| 1 5 | 15.31M | |
| 2 5 | 16.94M | |
| 7 10 | 16.94M | |
| 1 3 | 18.56M | |
| 3 5 | 18.56M | |
| 2 4 | 18.56M | |
| 9 11 | 18.56M | |
| 2 3 | 20.19M | |
| 3 4 | 20.19M | |
| 4 5 | 20.19M | |
| 9 10 | 20.19M | |
| - | 21.82M |
A.2.2 PCA Visualization
In Section 4.1 we show the effect of TBA on latent representations. In this section, we provide additional visualization using PCA for DiNO-S, DEiT-S, ViT-S, with different datasets and approximating both early and late blocks (see Figures 10, 11, 9, 7 and 8).
A.2.3 Image Classification
This section presents additional experiments that complement and extend those detailed in Section 4.2. Datasets and models are the ones detailed in Tables 7 and 6.
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
@@ A.2.3 Image Classification
| Accuracy | |||||||
| 112 | -127.02M | ||||||
| 04 | -46.19M | ||||||
| 15 | -46.19M | ||||||
| 25 | -34.64M | ||||||
| 710 | -34.64M | ||||||
| 02, 34 | -34.64M | ||||||
| 12, 34 | -23.09M | ||||||
| 13 | -23.09M | ||||||
| 24 | -23.09M | ||||||
| 35 | -23.09M | ||||||
| 57 | -23.09M | ||||||
| 01 | -11.55M | ||||||
| 12 | -11.55M | ||||||
| 23 | -11.55M | ||||||
| 34 | -11.55M | ||||||
| 45 | -11.55M | ||||||
| 56 | -11.55M | ||||||
| 67 | -11.55M | ||||||
| 78 | -11.55M | ||||||
| 89 | -11.55M | ||||||
| 910 | -11.55M | ||||||
| 1011 | -11.55M | ||||||
| 1112 | -11.55M | ||||||
| 1213 | -11.55M | ||||||
| 1314 | -11.55M | ||||||
| 1415 | -11.55M | ||||||
| 1516 | -11.55M | ||||||
| 1617 | -11.55M | ||||||
| 1718 | -11.55M | ||||||
| 1819 | -11.55M | ||||||
| 1920 | -11.55M | ||||||
| 2021 | -11.55M | ||||||
| 2122 | -11.55M | ||||||
| 2223 | -11.55M | ||||||
| original | 304.35M | ||||||
A.2.4 Transformation Generalization
In this section, we extend the analysis to evaluate the model’s ability to generalize transformations across datasets to different approximation blocks and models (i.e., DiNO-S).
| Accuracy | ||||||
| Encoder | Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| ViT-S | 2 3 | MNIST | ||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 4 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 3 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 8 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 11 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 7 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| Accuracy | ||||||
| Encoder | Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| DiNO-S | 2 3 | MNIST | ||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 4 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 3 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 8 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 11 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 7 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
A.2.5 Comparison
In this section, we provide additional results for the SkipAt baseline.
| Accuracy | |||||
| Skip | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| 1 5 | |||||
| 2 5 | |||||
| 7 10 | |||||
| 1 3 | |||||
| 3 5 | |||||
| 2 4 | |||||
| 8 10 | |||||
| 9 11 | |||||
| 0 1 | |||||
| 1 2 | |||||
| 2 3 | |||||
| 3 4 | |||||
| 4 5 | |||||
| 5 6 | |||||
| 6 7 | |||||
| 7 8 | |||||
| 8 9 | |||||
| 9 10 | |||||
| 10 11 | |||||
| original | |||||
@@ A.2.6 Analysis of misclassifications
In this section, we examine changes in per-class accuracy and misclassification patterns. As shown in Figure 12, models behave differently at block approximations. DiNO-S remains remarkably stable across blocks and classes, with the only degradation appearing for classes dog (when approximating blocks 10 or 11) and deer (for block 10 approximation). ViT-S shows a similar drop for class dog on its final block. Instead, the most noticeable hit occurs for class cat when the earlier blocks are approximated. For DEiT-S, several mid-to-late block approximations improve accuracy for various classes, whereas the very first block causes a clear relative decline in nearly every class. These observations suggest strategies like preferring late-block approximation for DEiT-S, or reserving extra samples for the linear transformation in order to recover the accuracy of difficult classes for the model.
In order to further investigate how the predictions change while approximating blocks, we plot the difference in the normalized confusion matrix before and after the approximation. In Figure 13, we show the delta confusion matrix for DEiT-S on CIFAR-100C. Also, here we can see how approximating the very first block makes the model puzzling and lose per-class accuracy (i.e., negative delta along the diagonal). On the other hand, approximating the last block acts as a regularizer, resulting in an overall gain in the per-class accuracy and, as a consequence, fewer misclassifications (negative deltas off-diagonal). This supports results shown in Figures 12 and 1.
Additionally, Figure 14 shows representative CIFAR-10 images that become misclassified after approximating a block of ViT-S. The patterns we observe mirror the trends in Figures 12 and 13: when approximating earlier blocks, we observe many images belonging to class cat to be misclassified. Instead, when approximating later blocks, we observe images of the class dog to be misclassified. Together, these qualitative examples show that understanding these block-specific vulnerabilities allows us to steer the approximation procedure, informing choices about which blocks to approximate based on the observed impact on the final model’s class-wise performance.

### v6

@@ 6 Conclusion
In this work, we first analyze the emergence of consistent block-wise representation similarities within pretrained foundation models and then propose a method to leverage these similarities to obtain smaller and more efficient yet performant models. To this end, we propose Transformer Optimization using Adaptive and Simple Transformations (TOAST), a novel method for easily approximate entire transformer blocks using a simple transformation, without requiring additional training or fine-tuning. Our extensive empirical evaluations across multiple pretrained vision models and datasets validate that TOAST significantly reduces model parameters while maintaining, and sometimes even improving, downstream task performance. Furthermore, TOAST’s straightforward linear approach often achieves better results than existing strategies like block skipping, and can be more effective than complex, trained approximations. TOAST thus offers a practical and efficient method for streamlining foundation models, making them more computationally accessible, and towards deployment in resource-constrained scenarios such as on-device settings.
@@ A.1 Implementation details
This section details the experiments conducted in Section˜4, providing information to reproduce them. Additionally, we provide the code as a zip file in the supplementary material.
@@ A.1.1 Models and Datasets
Table˜5 contains the full list of the pretrained models, while Table˜6 contains dataset information.
@@ A.1.4 Computational resources
A.1.5 Efficiency metrics
We evaluated model efficiency using two primary metrics. GFLOPs were used to measure the hardware-independent theoretical complexity of a single forward pass, calculated using the fvcore analysis library. Throughput, measured in samples per second, was used to quantify the practical, hardware-dependent inference speed. This was benchmarked by averaging the wall-clock time over numerous iterations on a single NVIDIA H100 GPU with a consistent batch size of 256.
@@ A.2 Additional Experiments
This section presents supplementary experiments to extend those detailed in Section˜4.
A.2.1 Latent Analysis
This section extend the analysis conducted in Section˜4.1, to analyze block-wise internal similarities, to additional models of different dimensionality: ViT-T, ViT-S, ViT-B and DiNO-S. Additionally, we provide visualization using PCA for DiNO-S, DEiT-S, ViT-S, with different datasets and approximating both early and late blocks (see Figures˜10, 11, 9, 7 and 8).
A.2.2 Image Classification
This section presents additional experiments that complement and extend those detailed in Section˜4.2. Datasets and models are the ones detailed in Tables˜6 and 5.
@@ A.2.2 Image Classification
A.2.3 Zero-shot Image Classification
To further assess the effectiveness of our approach, we evaluate TOAST in a zero-shot image classification setting. This evaluation utilizes the OpenCLIP-ViT-B model (Radford et al., 2021), which was pretrained on LAION-2B Schuhmann et al. (2022), with ImageNet1k serving as the downstream evaluation dataset.
| Params. | Approx. | Accuracy | |
| -6.49M | 0 1 | 57.93 | - |
| 1 2 | 64.20 | - | |
| 2 3 | 66.35 | -5.51% | |
| 3 4 | 64.65 | - | |
| 4 5 | 64.86 | - | |
| 5 6 | 58.05 | - | |
| 6 7 | 61.56 | - | |
| 7 8 | 58.53 | - | |
| 8 9 | 52.32 | - | |
| 9 10 | 59.21 | - | |
| 10 11 | 22.64 | - | |
| 149.07M | original | 70.21 | – |
The analysis is conducted only on the base version, as larger versions (e.g., OpenCLIP-ViT-L or OpenCLIP-ViT-H) contain too many parameters and are thus beyond the scope of this paper. As in previous experiments, the model remains frozen, and block approximations are computed using a shared linear transformation applied across all tokens, based on a subset of 3,000 training samples. Importantly, we apply these approximations only to the vision encoder, leaving the text encoder unchanged. We follow the standard ImageNet1k prompt templates. The results in Table˜11 lead to the conclusion that the impact on zero-shot accuracy is highly dependent on the targeted block’s position. The choice of which blocks to approximate is therefore crucial. For instance, approximating an early block (e.g., 1 2 or 2 3) results in a modest accuracy drop (e.g., 5.51%), yielding a competitive model with fewer parameters. In contrast, approximating the final block (i.e., 10 11) causes a catastrophic performance collapse of 67.75%. Meaning that, for OpenCLIP-ViT-B, later layers in the vision encoder appear to capture uniquely critical information for zero-shot generalization that cannot be effectively replicated by earlier ones. To the best of our knowledge, this work is the first to investigate training-free model size reduction in this challenging setting.
A.2.4 Evaluation with original classification heads
| Encoder | Approximation | Accuracy |
DEiT-S |
78.87 | |
| 78.72 | ||
| 78.39 | ||
| 75.29 | ||
| 74.36 | ||
| 72.34 | ||
| 69.30 | ||
| 66.12 | ||
| original | 79.66 | |
ViT-S |
78.25 | |
| 74.50 | ||
| 71.23 | ||
| 70.46 | ||
| 66.18 | ||
| 59.69 | ||
| 58.36 | ||
| 23.85 | ||
| original | 79.86 |
As mentioned in the main paper, our primary evaluation involves training a new linear classifier on top of the frozen model backbone to simulate a realistic transfer learning scenario. However, the original papers for DEiT-S Touvron et al. (2021) and ViT-S Beyer et al. (2022) report performance using the classification head that was part of the original pre-training.
To confirm that our conclusions are robust and not an artifact of our evaluation protocol, we conducted an additional set of experiments using the official, pre-trained classification heads from the original model checkpoints. In this setup, we do not train a new classifier; we simply evaluate the accuracy of the frozen, approximated models using their original heads.
The results, presented in Table˜12, are fully consistent with the main conclusions of our paper. They confirm that our block approximation method provides a favorable accuracy-efficiency trade-off, even when evaluated with the original model heads. The relative drop in accuracy when approximating different layers follows the same patterns observed in our primary experiments, reinforcing the validity of our approach.
A.2.5 Computational efficiency vs. accuracy
To quantify the effectiveness of different approximation methods, we analyze the trade-off between downstream accuracy and computational cost. Figure˜12 presents this analysis on a DiNO-B model using both CIFAR-100F and ImageNet1k against three standard efficiency metrics: parameter count, GFLOPs, and inference throughput. Across all metrics, the proposed linear translator (green) establishes a more favorable Pareto frontier compared to the baseline identity-based approach (blue). This indicates that for any given efficiency budget (e.g., a specific GFLOPs target), the linear translator consistently yields a model with higher accuracy.
@@ A.2.6 Analysis of misclassifications
In this section, we examine changes in per-class accuracy and misclassification patterns. As shown in Figure˜13, models behave differently at block approximations. DiNO-S remains remarkably stable across blocks and classes, with the only degradation appearing for classes dog (when approximating blocks 10 or 11) and deer (for block 10 approximation). ViT-S shows a similar drop for class dog on its final block. Instead, the most noticeable hit occurs for class cat when the earlier blocks are approximated. For DEiT-S, several mid‑to‑late block approximations improve accuracy for various classes, whereas the very first block causes a clear relative decline in nearly every class. These observations suggest strategies like preferring late‑block approximation for DEiT-S, or reserving extra samples for the linear transformation in order to recover the accuracy of difficult classes for the model.
In order to further investigate how the predictions change while approximating blocks, we plot the difference in the normalized confusion matrix before and after the approximation. In Figure˜14, we show the delta confusion matrix for DEiT-S on CIFAR-100C. Also, here we can see how approximating the very first block makes the model puzzling and lose per-class accuracy (i.e., negative delta along the diagonal). On the other hand, approximating the last block acts as a regularizer, resulting in an overall gain in the per-class accuracy and, as a consequence, fewer misclassifications (negative deltas off-diagonal). This supports results shown in Figures˜13 and 3.
Additionally, Figure˜15 shows representative CIFAR-10 images that become misclassified after approximating a block of ViT-S. The patterns we observe mirror the trends in Figures˜13 and 14: when approximating earlier blocks, we observe many images belonging to class cat to be misclassified. Instead, when approximating later blocks, we observe images of the class dog to be misclassified. Together, these qualitative examples show that understanding these block-specific vulnerabilities allows us to steer the approximation procedure, informing choices about which blocks to approximate based on the observed impact on the final model’s class-wise performance.